OpenStudy (anonymous):

Explain how to solve the equation 3t + 11 = –10. Show that your solution works.

OpenStudy (anonymous):

\[3t=21\] \[3t/3=21/3\] \[t=7\]

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (anonymous):

3t+11=-10 3t=-10-11 3t=-21 t=-21/3=-7

OpenStudy (anonymous):

^ -21!!